Present a complete process in computing the mean of the sampling distribution of the sample


means given the General Mathematics 20- item quiz result of 6 students; 12, 16, 18, 20, 10, 15.


Consider a sample size of 5.

1
Expert's answer
2021-11-17T02:40:22-0500

Solution:

Given result of 6 students; 12, 16, 18, 20, 10, 15.

Here, a sample size of 5, so n=5

We consider no repetition.

Samples are: (12, 16, 18, 20, 10), (16, 18, 20, 10, 15), (12, 18, 20, 10, 15), (12, 16, 20, 10, 15), (12, 16, 18,10, 15), (12, 16, 18, 20,15).

Their means are:

Mean of (12, 16, 18, 20, 10)"=(12+16+18+20+10)\/5=15.2"

Mean of (16, 18, 20, 10, 15)"=(16+18+20+10+15)\/5=15.8"

Mean of (12, 18, 20, 10, 15)"=(12+15+18+20+10)\/5=15"

Mean of (12, 16, 20, 10, 15) "=(12+16+15+20+10)\/5=14.6"

Mean of (12, 16, 18,10, 15) "=(12+16+18+15+10)\/5=14.2"

Mean of (12, 16, 18, 20,15)"=(12+16+18+20+15)\/5=16.2"

Now, mean of sampling distribution "=(15.2+15.8+15+14.6+14.2+16.2)/6=15\dfrac16"